OpenStudy (anonymous):

PLEASE HELP !! WILL GIVE MEDALL !! Which is a solution of the equation? 2x – y = 5 7 (2, –1) (1, 2) (3, –1)

OpenStudy (kc_kennylau):

what's the first option

OpenStudy (anonymous):

just 7

OpenStudy (anonymous):

(2, -1)

OpenStudy (kc_kennylau):

2x-y=5 x=(5+y)/2 y=2x-5

OpenStudy (anonymous):

thaaaanks alot (: @Nemo1223

OpenStudy (anonymous):

No Problem.